OpenStudy (anonymous):

Peter, John and Dan shared $1458 equally. Peter used part of his share to buy a bicycle and had $139 left. what was the cost of the bicycle??? please help i am totally lost

OpenStudy (anonymous):

WOULD TAKE 1458 /2=729.00 EACH THEN TAKE 139 FROM 729 AND WILL TELL YOU HOW MUCH HE SPENT FOR BIKE

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

BECAUSE TWO PEOPLE

OpenStudy (anonymous):

HOLD THREE PEOPLE

OpenStudy (anonymous):

if Peter, John, and Dan had to share 1458 equally you should divide 1458 by 3. 1458/3= 486 Each person had $486 to themselves so you would then subtract 139 486-139=347 the bicycle cost $347 dollars

OpenStudy (anonymous):

YEA

OpenStudy (anonymous):

SAME AS I SAID LOL BUT I READ IT TO FAST LOL

OpenStudy (anonymous):

SO INSTEAD OF 2 WAS 3 LOL